A balloon is rising vertically upward with a uniform


velocity 'u' as shown in Fig. 11.E22. At a certain


instant, a ball No. 1 is dropped from it, which takes a


4 sec to reach the ground. The moment this ball


strikes the ground, another ball is dropped from the


balloon and this 2nd ball takes 5 sec to reach the


ground. Find the velocity of the balloon 'u' and height


from which both the balls must have been dropped.

1
Expert's answer
2022-02-21T14:06:50-0500

"h_1=\frac{gt_1^2}2=80~m,"

"h_2=\frac{gt_2^2}2=125~m,"

"u=\frac{h_2-h_2}{t_1}=11. 25~\frac ms."